# Generalization or Memorization: Data Contamination and Trustworthy Evaluation for Large Language Models

### v1

@@ 3.1 Edit Distance Distribution
$\rho(d)=\frac{\sum\limits_{i=0}^{n}\sum\limits_{j=i+1}^{n}\mathbb{I}(\operatorname{ED}(s_{i},s_{j})=d)}{n*(n-1)/2},$ (2)
@@ 3.2 CDD for Data Contamination Detection
$\rho^{\prime}(d)=\frac{\sum\limits_{i=0}^{n}\mathbb{I}(\operatorname{ED}(s_{i},y)=d)}{n}.$ (3)
@@
$\rho^{*}(d)=\frac{\sum\limits_{i=0}^{n}\mathbb{I}(\operatorname{ED}(s_{i},s_{t=0})=d)}{n}.$ (4)
@@ 4 Experiment
In this section, we first introduce two datasets, DetCon and ComiEval, tailored for the tasks of data contamination detection and contamination mitigation evaluation, respectively (Section 4.1). We then evaluate the efficacy of CDD on the DetCon dataset (Section 4.2). Following this, we assess the performance of TED on the ComiEval dataset (Section 4.3). Finally, we demonstrate the application results of both CDD and TED in the real-world scenario (Section 4.4).
Table 1: Detailed statistics of simulating different data contamination scenes of LLMs.
Scenario Leaked Dataset Base LLMs Other Training Data Mixing Ratio Learning Rate Occurrences Contamination Form
@@
Approach DetCon (Code Generation) Avergae DetCon (Logical Reasoning) Avergae
@@
Embedding ${}^{\dagger}$ 0.524 0.569 0.571 0.554 0.592 0.645 0.668 0.635
${}^{\dagger}$ 0.513 0.593 0.491 0.532 0.497 0.664 0.699 0.620
${}^{\dagger}$ 0.563 0.524 0.565 0.550 0.527 0.677 0.698 0.634
LLM ${}^{\dagger}$ 0.535 0.578 - 0.556 0.509 0.433 - 0.471
CDD 0.715 0.694 0.761 0.724 ( normal-$\uparrow$ 30.2%) 0.706 0.765 0.846 0.773 ( normal-$\uparrow$ 21.8%)
@@ 4.1 Dataset
To simulate different data contamination scenes of LLMs, we conduct extensive experiments that span two representative downstream scenarios, two contamination forms, four mixing numbers, three learning rates, and twenty-one occurrences (i.e., contamination degrees) on four different base LLMs around 7B. The detailed statistics can be found in Table 1. We employ LoRA Hu et al. (2022) to fine-tune the base models on these various settings On this basis, we construct the DetCon and ComiEval datasets.
@@
contains 2224 data contamination detection tasks, covering two scenarios (code generation and logical reasoning) and two contamination forms (original and variant), which need to detect whether a specific LLM has contamination on a particular data. We randomly select the data from the leaked dataset and the LLM from the settings in Table 1, where occurrence 0 refers to ‘uncontaminated’ and the others denote ‘contaminated’.
@@
Figure 4: The influence of hyper-parameters, where $\alpha$ and $\xi$ serve for CDD, $\gamma$ is used for TED, and we use the gray dashed line to represent the employed hyper-parameters.
@@ 4.2 Data Contamination Detection
Figure 6: The impact of data contamination with different settings and occurrences.
@@
As presented in Table 3, compared with other contamination detection approaches, CDD attains SOTA performance in both code generation and logic reasoning scenarios. CDD exhibits steady improvements across the Accuracy, F1 Score, and AUC metrics, with the average relative improvement ranging between 21.8% and 30.2%. Moreover, the advantage of CDD is that it only requires the sampled texts of LLMs to detect data contamination, without the need for additional conditions in Table 2.
@@ 4.3 Contamination Mitigation Evaluation
We evaluate the effectiveness of TED for contamination mitigation in different learning rates, base LLMs, mixing ratios, contamination forms, and occurrences on ComiEval. We set the hyper-parameter $\gamma=2$ and use Pass@1 Chen et al. (2021) as the evaluation metric $\mathcal{E}$.
@@
TED can steadily mitigate the performance improvements across different settings and occurrences in data contamination scenes, as shown in Figure 6. Moreover, the advantage of TED is that the performance influence of TED on the uncontaminated model (i.e. 0 occurrences) is small and almost negligible. However, as contamination degrees continue to increase, the performance influence of TED becomes apparent in all of the different settings.
@@
As illustrated in Figure 4 (c), an increase in the hyperparameter $\gamma$ leads to a more pronounced suppression of performance improvements attributable to data contamination by TED. Meanwhile, it also marginally decreases the performance of the uncontaminated model.
@@ 4.4 Real-World Application
In real-world applications, we apply CDD and TED for ChatGPT and construct two new datasets to assist evidence: 1) CodeForces2305 comprises 90 of the easiest level programming problems collected from the CodeForces website since May 2023, which is after the most recent update deadline of ChatGPT’s training data, i.e., April 2023. 2) HumanEval_R is reconstructed on HumanEval, which replaces its function signature, translates its requirements into German, French, and Chinese, selects different public test cases from the work Dong et al. (2023a) to prompt, and remains the private test cases for testing. To enhance the detection precision, we set the hyper-parameters $\alpha$ to 0 and $\xi$ to a larger value of 0.2 for CDD. We keep $\gamma$ at the default value of 2 for TED.
Table 5: Data contamination detection and contamination mitigation evaluation for ChatGPT, where ‘pre’ and ‘post’ present ChatGPT’s APIs with fixed versions ‘0613’ and ’1106’ respectively, Avg. Peak means the average of the peakedness of output distribution computed via Eq. 5, and Leak Ratio means the percentage of leaked tasks detected by CDD in the benchmark.
@@ 7 Limitations
Finally, in constructing our datasets, we assume that the base LLMs used do not suffer from data leakage on the selected benchmarks. However, in reality, these LLMs may have slight data contamination. To completely avoid this issue, it might be necessary to retrain an LLM from scratch on a training set known to be entirely free of test data. However, undertaking such a process would be prohibitively costly.
@@ Appendix A Details of Dataset Construction
    LLMs. For code generation tasks, we use CodeLlama-7B Rozière et al. (2023) and CodeGen-6.7B Nijkamp et al. (2023); for natural language processing tasks, we select Llama2-7B and Bloom-7B.
@@
    Training Data. Code generation tasks use the training data from StarCoder Li et al. (2023a), while natural language processing tasks use RedPajama Computer (2023).
Next, we construct the dataset DetCon for data contamination detection, starting with the construction of uncontaminated samples. We directly use the outputs generated by LLMs on the test data, representing uncontaminated data. Then, we construct contaminated samples by simulating different contamination scenes:
@@
By combining these four different scenes, we can construct a variety of composite data contamination scenes. For each piece of test data, we randomly select the results generated by LLMs under one of the contamination scenes as the contaminated samples. Following the previous works Jiang et al. (2023); Dong et al. (2023b, c); Li et al. (2024), in generating these samples, we also record the outputs of greedy search with a temperature parameter of 0 (1 sample) and 50 samples obtained by sampling with a temperature of 0.8.
This construction approach aims to comprehensively cover possible data contamination scenes, ensuring we can accurately assess the performance of LLMs in the face of different types and degrees of data contamination.

### v2

@@ Abstract
[…]
[ tikzsetting= fill=blueish, skipabove=0.33em, skipbelow=0.33em, linewidth=1pt, innerleftmargin=4pt, innerrightmargin=4pt, innertopmargin=2pt, innerbottommargin=2pt, linecolor=gray95, roundcorner=2pt, shadowsize=4pt, shadowcolor=gray95 ]answerbox
@@ 3.1 Edit Distance Distribution
$\rho(d)=\frac{\sum\limits_{i=1}^{n-1}\sum\limits_{j=i+1}^{n}\mathbb{I}(\operatorname{ED}(s_{i},s_{j})=d)}{n*(n-1)/2},$ (2)
@@ 3.2 CDD for Data Contamination Detection
$\rho^{\prime}(d)=\frac{\sum\limits_{i=1}^{n}\mathbb{I}(\operatorname{ED}(s_{i},y)=d)}{n}.$ (3)
@@
$\rho^{*}(d)=\frac{\sum\limits_{i=1}^{n}\mathbb{I}(\operatorname{ED}(s_{i},s_{t=0})=d)}{n}.$ (4)
@@ 4 Experiment
[…]
Table 1: Detailed statistics of simulating different data contamination scenarios of LLMs.
Domain Leaked Dataset Base LLMs Other Training Data Mixing Ratio Learning Rate Occurrences Contamination Form
@@
Approach DetCon (Code Generation) Average DetCon (Logical Reasoning) Average
@@
Embedding similarity 0.524 0.569 0.571 0.554 0.592 0.645 0.668 0.635
Perplexity 0.513 0.593 0.491 0.532 0.497 0.664 0.699 0.620
Min-kProb 0.563 0.524 0.565 0.550 0.527 0.677 0.698 0.634
LLM Decontaminator 0.535 0.578 - 0.556 0.509 0.433 - 0.471
CDD 0.715 0.694 0.761 0.724 ( $\uparrow$ 30.2%) 0.706 0.765 0.846 0.773 ( $\uparrow$ 21.8%)
@@ 4.1 Dataset
We simulate data contamination by training LLMs using benchmark data. To cover various scenarios of data contamination, we consider different settings, including two domain benchmarks leaked on four LLMs, two contamination form (i.e. original and variant leaked data), using three different learning rates during training, four mixing ratios of leaked data with other training data, and 21 degrees of contamination (i.e., occurrences). The detailed statistics can be found in Table 1. Due to the high cost of large-scale pre-training, we employ LoRA Hu et al. (2022) to fine-tune the base models on these various settings. On this basis, we construct the DetCon and ComiEval datasets.
@@
contains 2224 data contamination detection tasks, covering two domains (code generation and logical reasoning) and two contamination forms (original and variant), which need to detect whether a specific LLM has contamination on a particular data. We randomly select the data from the leaked dataset and the LLM from the settings in Table 1, where occurrence 0 refers to ‘uncontaminated’ and the others denote ‘contaminated’.
@@
Figure 4: The influence of hyper-parameters, where $\alpha$ and $\xi$ serve for CDD, $\tau$ is used for TED, and we use the gray dashed line to represent the employed hyper-parameters.
@@ 4.2 Data Contamination Detection
Figure 6: The effect of our TED on model performance as the degree of data contamination increases under different settings. The legend displays the settings for specific data leakage scenarios.
@@
As presented in Table 3, compared with other contamination detection approaches, CDD attains SOTA performance in both code generation and logic reasoning domains. CDD exhibits steady improvements across the Accuracy, F1 Score, and AUC metrics, with the average relative improvement ranging between 21.8% and 30.2%. Moreover, the advantage of CDD is that it only requires the sampled texts of LLMs to detect data contamination, without the need for additional conditions in Table 2.
@@ 4.3 Contamination Mitigation Evaluation
We evaluate the effectiveness of TED for contamination mitigation in different learning rates, base LLMs, mixing ratios, contamination forms, and occurrences on ComiEval. We set the hyper-parameter $\tau=2$ and use Pass@1 Chen et al. (2021) as the evaluation metric $\mathcal{E}$.
@@
TED can steadily mitigate the performance improvements across different settings and occurrences in data contamination scenarios, as shown in Figure 6. Moreover, the advantage of TED is that the performance influence of TED on the uncontaminated model (i.e. 0 occurrences) is small and almost negligible. However, as contamination degrees continue to increase, the performance influence of TED becomes apparent in all of the different settings.
@@
As illustrated in Figure 4 (c), an increase in the hyperparameter $\tau$ for TED leads to a more pronounced suppression of performance improvements attributable to data contamination. Meanwhile, it also marginally decreases the performance of the uncontaminated model.
@@ 4.4 Real-World Application
In real-world applications, we apply CDD and TED for ChatGPT and construct two new datasets to assist evidence: 1) CodeForces2305 comprises 90 of the easiest level programming problems collected from the CodeForces website since May 2023, which is after the most recent update deadline of ChatGPT’s training data, i.e., April 2023. 2) HumanEval_R is reconstructed on HumanEval, which replaces its function signature, translates its requirements into German, French, and Chinese, selects different public test cases from the work Dong et al. (2023a) to prompt, and remains the private test cases for testing. To enhance the detection precision, we set the hyper-parameters $\alpha$ to 0 and $\xi$ to a larger value of 0.2 for CDD. We keep $\tau$ at the default value of 2 for TED.
Table 5: Data contamination detection and contamination mitigation evaluation for ChatGPT, where ‘pre’ and ‘post’ present ChatGPT’s APIs with fixed versions ‘0613’ and ’1106’ respectively, Avg. Peak means the average of the peakedness of output distribution computed via Eq. 5, and CR means the ratio of contaminated tasks detected by CDD in the benchmark.
@@ 7 Limitations
Finally, in constructing our datasets, we assume that the four base LLMs used do not suffer from data contamination on the selected benchmarks. However, in reality, these LLMs may have slight data contamination. To completely avoid this issue, it might be necessary to retrain an LLM from scratch on a training set known to be entirely free of test data. However, undertaking such a process would be prohibitively costly.
@@ Appendix A Details of Dataset Construction
In this section, we further describes the different data contamination scenarios, as well as how we collect and process data from these scenarios to construct the dataset.
@@
    LLMs. For code generation tasks, we use CodeLlama-7B Rozière et al. (2023) and CodeGen-6.7B Nijkamp et al. (2023); for logical reasoning tasks, we select Llama2-7B Touvron et al. (2023) and Bloom-7B Scao et al. (2022).
@@
    Training Data. Code generation tasks use the training data from StarCoder Li et al. (2023a), while logical reasoning tasks use RedPajama Computer (2023).
Next, we construct the dataset DetCon for data contamination detection, starting with the construction of uncontaminated samples. We directly use the outputs generated by LLMs on the test data, representing uncontaminated data. Then, we construct contaminated samples by simulating different contamination scenarios:
@@
By combining these four different scenarios, we can construct a variety of composite data contamination scenarios. For each piece of test data, we randomly select the results generated by LLMs under one of the contamination scenarios as the contaminated samples. Following the previous works Jiang et al. (2023); Dong et al. (2023b, c); Li et al. (2024), in generating these samples, we also record the outputs of greedy search with a temperature parameter of 0 (1 sample) and 50 samples obtained by sampling with a temperature of 0.8.
This construction approach aims to comprehensively cover possible data contamination scenarios, ensuring we can accurately assess the performance of LLMs in the face of different types and degrees of data contamination.